# Solving Harder Reinforcement Learning Problems

### v1

@@ 5.2 RL on Gemma-2
We demonstrate promising RL results with the 2B variant of Gemma-2 [31] on real-world problems from MATH [17] and the easy public subset of ARC-AGI-1 [7]. The latter is a challenging reasoning task-set even for much larger state-of-the-art models than Gemma-2.
@@
We use a v5litepod-128 [14] which needs around 4 hours per 1000 training steps. An RL run for LLM training [29] involves sampling a fixed $n$ number of completions $\{x_{i}\}_{i=1}^{n}$ for a given prompt at a given training step. For our experiments, we set $n=16$. The rewards are computed for every completion using a reward function $g(\cdot)$. We transform these rewards $\{g(x_{i})\}_{i=1}^{n}$ using our unbiased estimator $s^{(\mathrm{loo}-1)}$ of (33), which we favour due to Figure 8, and which we refer to as PKPO. We repeat the training for a selection of $k^{\mathrm{opt}}$, thus optimizing a different $\mathrm{pass@k^{opt}}$ each time. Since $k^{\mathrm{opt}}=1$ leads to no reward transformation, this is our baseline (although we use basic LOO mean centering of Equation 26, without which the training diverges). For each run, we measure $\mathrm{pass@k^{eval}}$ for every $k^{\mathrm{eval}}$$\in\{1,2,4,8,12,16\}$ at each step. Additionally, we also track model entropy and cumulative solve rate during training. The latter is defined as the fraction of tasks from the task-set for which the model has sampled a correct solution at least once.
(a) Cumulative solve rate ($N=12000$ problems)
(b) Effect of changing $k$ on entropy during training
Figure 2: (a): Increasing $k^{\mathrm{opt}}$ in PKPO training solves more problems during Gemma-2 RL. (b): A higher $k^{\mathrm{opt}}$ makes the model learn to have higher entropy during RL. Thus, by optimizing for $\mathrm{pass@k}$  with $k>1$ instead of $\mathrm{pass@1}$, the model tends to have higher entropy leading to better exploration and solving more problems. Note that the size of one epoch which is 750 steps is evident in (a), where we see the slope decrease at each epoch boundary.
@@ 5.2.1 Choosing $k^{\mathrm{opt}}$  selectively optimizes $\mathrm{pass@k^{eval}}$  and solves more tasks
In Figure 3, we compare $\mathrm{pass@k^{eval}}$ across our runs ($k^{\mathrm{opt}}$$\in\{1,4,8\}$) for various $k^{\mathrm{eval}}$. We find the best $\mathrm{pass@k^{eval}}$ when $k^{\mathrm{opt}}=k^{\mathrm{eval}}$ (or $k^{\mathrm{opt}}$ is closest to $k^{\mathrm{eval}}$ among available $k^{\mathrm{opt}}$). Non-transformed rewards optimize $\mathrm{pass@1}$, leading to sub-optimal $\mathrm{pass@k^{eval}}$ for $k^{\mathrm{eval}}$$\neq 1$, and the deficit worsens as $k^{\mathrm{eval}}$ increases. Thus, our experiments also demonstrate that setting $k^{\mathrm{opt}}$$\coloneqq$ $k^{\mathrm{eval}}$ in our transformation suffices to optimize $\mathrm{pass@k^{eval}}$ for a $k^{\mathrm{eval}}$$\leq n$. This generalizes the already powerful result of [32] by alleviating the coupling that forces to either optimize $\mathrm{pass@n}$ or $\mathrm{pass@1}$. In other words, since RL training of LLMs typically samples a large batch ($n\gg 1$), not using our transformation leaves significant $\mathrm{pass@k}$ performance on the table for modest $k$.
@@
As $k^{\mathrm{opt}}\longrightarrow n$, the variance of our estimator increases as there are fewer subsets in (33) (see Figure 8). We presume this is why 1) gains of $k^{\mathrm{opt}}=8$ over $k^{\mathrm{opt}}=4$ are more prominent when $k^{\mathrm{eval}}$$\in\{12,16\}$ than when $k^{\mathrm{eval}}$ $=8$. That is, when $k^{\mathrm{eval}}$ is further away from $k^{\mathrm{opt}}\in\{4,8\}$ than when it is closer, and 2) the special case $n=k^{\mathrm{opt}}$ of [32] struggles to optimize the $\mathrm{pass@n}$.
(a) $\mathrm{pass}@1$ ($k^{\mathrm{eval}}=1$)
(b) $\mathrm{pass}@2$ ($k^{\mathrm{eval}}=2$)
(c) $\mathrm{pass}@4$ ($k^{\mathrm{eval}}=4$)
(d) $\mathrm{pass}@8$ ($k^{\mathrm{eval}}=8$)
(e) $\mathrm{pass}@12$ ($k^{\mathrm{eval}}=12$)
(f) $\mathrm{pass}@16$ ($k^{\mathrm{eval}}=16$)
Figure 3: Effect of $k^{\mathrm{opt}}$ (used in our PKPO training) on the rolling $\mathrm{pass@k^{eval}}$ in Gemma-2 RL. Setting $k^{\mathrm{opt}}=k^{\mathrm{eval}}$ usually achieves the best $\mathrm{pass@k^{eval}}$. Prior work [32] (which is equivalent to the specific case of $k^{\mathrm{opt}}=n=16$ in our notation) is also shown for comparison, and suffers here presumably due to the larger estimator variance and unreliable gradient (see also Figure 8).
5.2.2 Improving $\mathrm{pass@k^{eval}}$ without sacrificing $\mathrm{pass}@1$
@@
Figure 4 demonstrates that as PKPO can use any arbitrary $k^{\mathrm{opt}}\leq n$, this allows varying $k^{\mathrm{opt}}$ over the course of training to good affect. We show a simple annealing procedure which starts training with a high $k^{\mathrm{opt}}=8$ and reduces it to $k^{\mathrm{opt}}=1$ after 1500 steps. This trains the model to initially prioritize exploration (optimize $\mathrm{pass@k}$) and then consolidate the single-sample policy (optimize $\mathrm{pass@1}$). This switch is apparent in Figure 4(a), at step 1500 where the slope of $k^{\mathrm{annealed}}$ changes. While traditional methods like [32] suffer from a trade-off between $\mathrm{pass@k}$ and $\mathrm{pass@1}$, we get a final model which has higher $\mathrm{pass@k^{eval}}$ for all $k^{\mathrm{eval}}$$>1$ with no sacrifice in $\mathrm{pass@1}$.
@@
Figure 4: Annealing $k^{\mathrm{opt}}$ during PKPO training improves $\mathrm{pass}@k^{\mathrm{eval}}$ without sacrificing $\mathrm{pass}@1$. For $k^{\mathrm{annealed}}$ , we train with $k^{\mathrm{opt}}=8$ up to step $1500$ and $k^{\mathrm{opt}}=1$ thereafter.
5.2.3 For hard problems, transforming the rewards is essential for learning
Figure 5 shows the limitation of traditional $\mathrm{pass@1}$ optimization through RL on an especially challenging task-set. We use the easy subset of ARC-AGI-1 [7]. We observe that conventional $\mathrm{pass@1}$ optimization stalls. However, our $\mathrm{pass@k}$ approach unblocks learning, and results in higher $\mathrm{pass@k^{eval}}$ across all $k^{\mathrm{eval}}$ including $k^{\mathrm{eval}}$$=1$. Furthermore, we see higher $k^{\mathrm{opt}}$ leads to more effective and faster learning. This is likely because a harder task-set makes the benefit of prioritizing joint utility over individual sample utility more prominent.
(a) Cumulative solve rate
(b) $\mathrm{pass}@1$ ($k^{\mathrm{eval}}=1$)
(c) $\mathrm{pass}@16$ ($k^{\mathrm{eval}}=16$)
@@ 6 Conclusions and Outlook
In RL training with multiple independent samples per task, optimizing the $\mathrm{pass@k}$ maximizes the expectation of the best reward in the set of samples, rather than the average one. This retains model output diversity leading to the eventual solution of more problems, and so to stronger policies. We provide drop-in replacements for more traditional RL reward transformations that robustly and efficiently optimize the $\mathrm{pass@k}$. This work can be extended in various ways, such as to other inference-time search algorithms, and to more sophisticated baseline techniques.
Appendix A Additional Proofs
A.1 Proof of Theorem 2
@@
Due to the second condition, $m_{i}$ therefore equals one of two values which we denote by $m^{(1)}$ and $m^{(0)}$, depending on whether $f_{i}=1$ or $f_{i}=0$, respectively.
If $f_{i}=1$ then all subsets that include $i$ also include at least one correct element ($i$ itself), so that $m^{(1)}$ is just the number of subsets of size $k$ of $\{1,2,\dots,n\}$ that include $i$, which equals the number of subsets of size $k-1$ of $\{1,2,\dots,n-1\}$:
@@
If $f_{i}=0$ then we assume w.l.o.g. that $i=n$, so that $m^{(0)}$ is the number of subsets of size $k-1$ of $\{1,2,\dots,n-1\}$ with at least one correct element,
@@ A.2 Proof of Theorem 5
Because of the structure of $M$ we have that
@@ 45 N = len(g)
Listing 1: Python implementations of our reward transformations. Functions with names that begin with an underscore are helpers, while the remaining four functions rho, s, sloo and sloo_minus_one implement $\rho^{(g)}$, $s_{i}$, $s_{i}^{(\mathrm{loo})}$ and $s_{i}^{(\mathrm{loo}-1)}$, respectively. For simplicity this implementation costs $\mathcal{O}(nk+n\log n)$ — reducing this to $\mathcal{O}(k+n\log n)$ would require optimizing _deltas and _m_diagonal.

### v2

@@ 5.2 RL on Open Source LLMs
We demonstrate promising RL results with the 2B and 9B parameter variants of Gemma2 [38] and the 8B parameter variant of Llama3.1 on real-world problems in MATH [20], code generation [4] [12], and the easy public subset of ARC-AGI-1 [8]. The latter is a challenging reasoning task-set even for state-of-the-art models much larger than Gemma2.
@@
For Gemma2-2B we use a v5litepod-128 [17] which needs around 4 hours per 1000 training steps. Each RL training run [36] involves sampling a fixed $n$ number of completions $\{x_{i}\}_{i=1}^{n}$ for a given prompt at a given training step. For our experiments, we set $n=16$. The rewards are computed for every completion using a reward function $g(\cdot)$. We transform these rewards $\{g(x_{i})\}_{i=1}^{n}$ using our unbiased estimator $s^{(\mathrm{loo}-1)}$ of (33), which we favour due to Figure 4, and which we refer to as PKPO. We repeat the training for a selection of $k^{\mathrm{opt}}$, thus optimizing a different $\mathrm{pass@k^{opt}}$ each time. Since $k^{\mathrm{opt}}=1$ leads to no reward transformation, this is our baseline (although we use basic LOO mean centering of Equation 26, without which the training diverges). For each run, we measure $\mathrm{pass@k^{eval}}$ for every $k^{\mathrm{eval}}$$\in\{1,2,4,8,12,16\}$ at each step. Additionally, we also track model entropy and cumulative solve rate during training. The latter is defined as the fraction of tasks from the task-set for which the model has sampled a correct solution at least once; this is a critical metric that reflects the success of the model’s exploration and measures its ability to find novel solutions.
Table 5: Results for Gemma2-9B on the ARC-AGI-1 benchmark.
Gemma2-9B Cumulative Solve Rate pass@1 pass@16
k_opt=1 12.00 $\pm$ 04.33 02.00 $\pm$ 01.69 08.18 $\pm$ 04.00
k_opt=4 82.33 $\pm$ 04.14 22.00 $\pm$ 02.00 38.18 $\pm$ 04.67
k_opt=8 84.14 $\pm$ 04.67 26.67 $\pm$ 02.50 44.50 $\pm$ 04.33
[39] 22.00 $\pm$ 04.44 06.00 $\pm$ 02.67 10.16 $\pm$ 04.57
EntropyReg 24.67 $\pm$ 04.50 04.00 $\pm$ 02.33 08.89 $\pm$ 04.89
Table 6: Results for Llama3.1-8B on the ARC-AGI-1 benchmark.
Llama3.1-8B Cumulative Solve Rate pass@1 pass@16
k_opt=1 22.00 $\pm$ 04.18 03.33 $\pm$ 02.00 08.00 $\pm$ 04.50
k_opt=4 87.17 $\pm$ 04.14 24.33 $\pm$ 02.33 42.00 $\pm$ 04.16
k_opt=8 88.89 $\pm$ 04.33 29.67 $\pm$ 02.67 43.13 $\pm$ 04.67
[39] 36.00 $\pm$ 02.50 08.00 $\pm$ 04.00 18.00 $\pm$ 04.89
EntropyReg 28.00 $\pm$ 04.44 08.00 $\pm$ 02.50 14.67 $\pm$ 04.44
Entropy regularization baseline In addition to our PKPO and the special case thereof of [39], we also add the entropy regularization baseline, which is PPO with an additional entropy term in the objective. We give this baseline an arguably unfair advantage by performing a small sweep over the values $0.001,0.005,0.01,0.05,0.1$ for the entropy_coefficient for each (model, benchmark) pair and only report the best result as EntropyReg.
@@ 5.2.1 Choosing $k^{\mathrm{opt}}$ selectively optimizes $\mathrm{pass@k^{eval}}$ and solves more tasks
In Figure 7, we compare $\mathrm{pass@k^{eval}}$ across our runs ($k^{\mathrm{opt}}$ $\in\{1,4,8\}$) for various $k^{\mathrm{eval}}$. We find the best $\mathrm{pass@k^{eval}}$ when $k^{\mathrm{opt}}=k^{\mathrm{eval}}$ (or $k^{\mathrm{opt}}$ is closest to $k^{\mathrm{eval}}$ among available $k^{\mathrm{opt}}$). Non-transformed rewards optimize $\mathrm{pass@1}$, leading to sub-optimal $\mathrm{pass@k^{eval}}$ for $k^{\mathrm{eval}}$$\neq 1$, and the deficit worsens as $k^{\mathrm{eval}}$ increases. Thus, our experiments also demonstrate that setting $k^{\mathrm{opt}}\coloneqq$ $k^{\mathrm{eval}}$ in our transformation suffices to optimize $\mathrm{pass@k^{eval}}$ for a $k^{\mathrm{eval}}$$\leq n$. This generalizes the already powerful result of [39] by alleviating the coupling that restricts to optimizing either $\mathrm{pass@n}$ or $\mathrm{pass@1}$. In other words, since RL training of LLMs typically samples a large batch ($n\gg 1$), failing to use our transformation results in sub-optimal $\mathrm{pass@k}$ performance, especially for modest values of $k$.
@@
As $k^{\mathrm{opt}}\longrightarrow n$, the variance of our estimator increases as there are fewer subsets in (33) (see Figure 4). We presume this is why 1) gains of $k^{\mathrm{opt}}=8$ over $k^{\mathrm{opt}}=4$ are more prominent when $k^{\mathrm{eval}}$ $\in\{12,16\}$ than when $k^{\mathrm{eval}}$ =8=8. That is, when $k^{\mathrm{eval}}$ is further away from $k^{\mathrm{opt}}\in\{4,8\}$ than when it is closer, and 2) the special case $n=k^{\mathrm{opt}}$ of [39] struggles to optimize the $\mathrm{pass@n}$.
5.2.2 PKPO robustly improves $\mathrm{pass@k}$ on held out evaluations
Tables 1-4 above (and Tables 7-8 in the appendix) present performance on held-out sets for two tasks. We report the mean and standard error based on three runs with different random seeds. For math, we train on the train split and evaluate on the test split of Hendrycks MATH [20]. To evaluate coding, we use MBPP [4] for training and evaluate on HumanEval [12]. MBPP has multiple unit tests per problem and hence we use this not only as a proxy for additional benchmarks but also to showcase our handling of a continuous reward function (% unit tests passed).
5.2.3 Improving $\mathrm{pass@k}$ without sacrificing $\mathrm{pass}@1$
@@
Figure 3 demonstrates that as PKPO can use any arbitrary $k^{\mathrm{opt}}\leq n$, this allows varying $k^{\mathrm{opt}}$ over the course of training to good effect. We show a simple annealing procedure which starts training with a high $k^{\mathrm{opt}}=8$ and reduces it to $k^{\mathrm{opt}}=1$ after 1500 steps. This trains the model to initially prioritize exploration (optimize $\mathrm{pass@k}$) and then consolidate the single-sample policy (optimize $\mathrm{pass@1}$). This switch is apparent in Figure 3(a), at step 1500 where the slope of $k^{\mathrm{annealed}}$ changes. While traditional methods like [39] suffer from a trade-off between $\mathrm{pass@k}$ and $\mathrm{pass@1}$, we get a final model which has higher $\mathrm{pass@k^{eval}}$ for all $k^{\mathrm{eval}}$>1>1 with no sacrifice in $\mathrm{pass@1}$.
5.2.4 PKPO is essential for learning on hard problems
Figure 8 shows the limitation of traditional $\mathrm{pass@1}$ optimization through RL on an especially challenging task-set. We use the easy subset of ARC-AGI-1 [8]. We observe that conventional $\mathrm{pass@1}$ optimization stalls. However, our $\mathrm{pass@k}$ approach unblocks learning, and results in higher $\mathrm{pass@k^{eval}}$ across all $k^{\mathrm{eval}}$ including $k^{\mathrm{eval}}$=1=1. Furthermore, we see higher $k^{\mathrm{opt}}$ leads to more effective and faster learning. This is likely because the benefits of prioritizing joint utility over individual sample utility are more prominent on a harder task-set.
@@ 5.2.4 PKPO is essential for learning on hard problems
Figure 3: Annealing $k^{\mathrm{opt}}$ during PKPO training improves $\mathrm{pass}@k^{\mathrm{eval}}$ without sacrificing $\mathrm{pass}@1$. For $k^{\mathrm{annealed}}$ , we train with $k^{\mathrm{opt}}=8$ up to step 15001500 and $k^{\mathrm{opt}}=1$ thereafter.
Tables 5 and 6 show more extensive experiments on ARC-AGI-1. We make an 80:20 train:test split of the same easy subset as before and report the cumulative solve rate on the train set and $\mathrm{pass@k}$ rate on the test set. We train to saturation (no change in cumulative rate for 1k steps), and again use three random restarts to provide standard errors. By encouraging exploration in a direct and stable manner, our method unblocks learning unlike other methods. Entropy Regularization does indeed sacrifice pass@1 and slightly improves $\mathrm{pass@k}$ by promoting exploration, but it is hard to tune, and is significantly outperformed by our method. Moreover, it has no explicit way to optimize for a specific k_eval. [39] targets the same objective as PKPO, but couples the minibatch size to $k$ and thereby incurs higher variance than PKPO with $k<n$.
@@ 6 Conclusions and Outlook
In RL training with multiple independent samples per task, optimizing the $\mathrm{pass@k}$ maximizes the expectation of the best reward in the set of samples, rather than the average one. This preserves model output diversity, which leads to solving more problems and ultimately yields stronger policies. We provide drop-in replacements for more traditional RL reward transformations that robustly and efficiently optimize the $\mathrm{pass@k}$. This work can be extended in various ways, such as to other inference-time search algorithms, and to more sophisticated baseline techniques.
Appendix A Additional Theoretical Statements and Proofs
A.1 Statement and proof that $n\geq k$ samples are required to unbiasedly estimate $\mathrm{pass@k}$
This result is a direct consequence of a well-known theorem concerning the unbiased estimability of parametric functions for the Bernoulli distribution.
Theorem 8 (Kolmogorov [23]).
Let $Y_{1},\dots,Y_{n}$ be i.i.d. Bernoulli random variables with success probability $p\in[0,1]$. A function $\rho(p)$ is unbiasedly estimable from this sample if and only if it can be expressed as a polynomial in $p$ of degree at most $n$.
A sketch of a proof of Theorem 8 can be found in Lehmann and Casella [26].
Corollary 2.
Given a sequence of $n$ i.i.d. model samples $x_{1},x_{2},\dots,x_{n}$, the $\mathrm{pass@k}$ is unbiasedly estimable if and only if $n\geq k$.
Proof.
Let $p=\mathbb{P}\big[[f(x_{i})=1]\big]$ be the probability that any one sample is correct. The $\mathrm{pass@k}$ is the complement of the probability that all of $k$ samples are incorrect, $1-(1-p)^{k}$. Because this is a polynomial in $p$ of order $k$, the result follows immediately from Theorem 8. ∎
A.2 Characterization of the Variance
Our proof of Theorem 1 reveals that the classic Codex formula $\rho(n,c,k)$ is a $U$-statistic (as is our generalization $\rho^{(g)}(n,c,k)$). The concept of $U$-statistics was introduced by Hoeffding, who gave
Theorem 9 (Hoeffding [21]).
Let $X_{1},\ldots,X_{n}$ be independent and identically distributed random variables. Let $h(x_{1},\ldots,x_{k})$ be a symmetric “kernel” with $\mathbb{E}[h(X_{1},\ldots,X_{k})^{2}]<\infty$. Define the parameter $\mu=\mathbb{E}[h(X_{1},\ldots,X_{k})]$ and the $U$-statistic
$U_{n}=\binom{n}{k}^{-1}\sum_{1\leq i_{1}<\dots<i_{k}\leq n}h(X_{i_{1}},\ldots,X_{i_{k}}).$ (36)
Further, let $h_{1}(x)=\mathbb{E}[h(x,X_{2},\ldots,X_{k})]$ and define the variance $\zeta_{1}=\text{Var}(h_{1}(X_{1}))$.
If $\zeta_{1}>0$ (the non-degenerate case), then as $n\to\infty$, the $U$-statistic $U_{n}$ is asymptotically normal:
$\sqrt{n}(U_{n}-\mu)\xrightarrow{d}\mathcal{N}(0,k^{2}\zeta_{1}),$ (37)
where $\xrightarrow{d}$ denotes convergence in distribution.
For the $\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k}$ our kernel is $h(\cdot)\equiv\max(\cdot)$, and we have the following.
Corollary 3.
As $n\rightarrow\infty$ the variance of $\rho(n,c,k)$ decreases at a rate of $1/n$ with
$\mathbb{E}\left[(\rho(n,c,k)-\mathrm{max}_{\hskip-0.40974ptg\hskip-0.81949pt}\mathrm{@k})^{2}\right]=\frac{k^{2}}{n}\zeta_{1},$ (38)
where $\zeta_{1}$ is a constant defined in Theorem 9 that depends on the model $p(x|\theta)$ and the univariate reinforcement learning reward function $g$.
A.3 Proof of Theorem 2
Although Theorem 2 is a special case of Theorem 4, we include both because the following proof uses a different approach from that of the more general statement, and is arguably the easier of the two.
@@ A.3 Proof of Theorem 2
Due to the second condition, $m_{i}$ therefore equals one of two values, which we denote by $m^{(1)}$ and $m^{(0)}$, depending on whether $f_{i}=1$ or $f_{i}=0$, respectively.
If $f_{i}=1$ then all subsets that include $i$ also include at least one correct element (ii itself), so that $m^{(1)}$ is just the number of subsets of size $k$ of $\{1,2,\dots,n\}$ that include $i$, which equals the number of subsets of size k1k-1 of $\{1,2,\dots,n-1\}$:
@@
If $f_{i}=0$ then we assume w.l.o.g. that $i=n$, so that $m^{(0)}$ is the number of subsets of size k1k-1 of $\{1,2,\dots,n-1\}$ with at least one correct element,
@@ A.4 Proof of Theorem 5
Because of the structure of $M$, we have that
@@ 45 N = len(g)
Listing 1: Python reward batch transformations. Functions with names that begin with an underscore are helpers, while the remaining four functions rho, s, sloo and sloo_minus_one implement $\rho^{(g)}$, $s_{i}$, $s_{i}^{(\mathrm{loo})}$ and $s_{i}^{(\mathrm{loo}-1)}$, respectively. For simplicity this implementation costs $\mathcal{O}(nk+n\log n)$ — reducing this to $\mathcal{O}(k+n\log n)$ would require optimizing _deltas and _m_diagonal.
Appendix C Additional Figures
Figure 4: The variance of different estimators of the gradient of $\mathrm{max}_{\hskip-0.40251ptg\hskip-0.80504pt}\mathrm{@k}$ with $k=4$ for the one-dimensional problem depicted in Figure 1 at the location $x=1$. Each data-point is the sample variance of 10,000 independent unbiased gradient estimates (lower is better). The horizontal axis denotes the number of samples $n$ used to construct each of the 10,000 estimates. We compare the following methods:
all subsets baselined: $s^{(\mathrm{loo})}$ — our novel estimator of Equation 29 that analytically sums over all subsets of size $k$ of the $n$ samples with our unbiased baseline method that subtracts for each element $i$ the mean of the estimator over all subsets of size $k$ that do not include ii.
naive partitioned baselined — a naive transformation that sets all $k$ transformed rewards in a subset of $k$ samples equal to the largest raw reward in that subset. To extend this method to $n>k$ we partition the $n$ samples (for integer multiples of $k$) into disjoint subsets of size $k$ and average the estimated gradient obtained from each. Furthermore, as a simple variance reduction method, for each such set of $k$ samples we subtract the mean of the transformed rewards from the other sets of $k$ samples (thereby averaging over (nk)(n-k) samples and subtracting the result from the $k$ samples and repeating $n/k$ times in a leave-one-out fashion over the subsets of size $k$). If we were to randomly sample an increasing number of partitions of the samples and average over all of them, then intuitively the resulting estimator would approach the variance of $s^{(\mathrm{loo})}$, but this would be expensive and indeed the limiting case of considering all partitions is intractable for general $n$ and $k$. Our estimators have the key property of summing over all such partitions while nonetheless being efficient to compute.
loo minus one all subsets: $s^{(\mathrm{loo}-1)}$ — our novel estimator of Equation 33 that analytically sums over all subsets of size $k$ of the $n$ samples and uses all appropriate subsets of size k1k-1 to form the variance-reducing baseline that retains unbiasedness, thereby non-trivially generalizing [39] to all $n>k$ with strong variance reduction.